# MEDebiaser: A Human-AI Feedback System for Mitigating Bias in Multi-label Medical Image Classification

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06submissionid: 8158ccs: Human-centered computing  Human computer interaction (HCI)
@@ 3.1.3. Analysis
We transcribed the audio recordings into text scripts and corrected transcription errors. Using thematic analysis (Oktay, 2012), we conducted a comprehensive examination of the interview data. Initially, all authors read through the scripts to achieve a shared understanding, then proceeded with the coding process. During the initial coding phase, two authors segmented the text data into meaningful units and assigned labels to each segment. Subsequently, two other authors refined and adjusted these labels, uncovering additional patterns and themes. Finally, all authors shared and discussed the coding results, reaching a consensus to ensure consistency and reliability. Based on the coding results, we summarized the codebook (Table 2), which includes $4$ themes and $13$ labels, with definitions clarified based on the interviewees’ perspectives. Through this rigorous process, we addressed RQ1 and RQ2, summarizing the challenges and needs faced by physicians and engineers in MLMIC during their collaboration with AI.
@@ 3.2.1. Challenges in Multi-label Medical Image Classification
Physician’s Perspective. All physicians underscored the unique complexities of multi-label medical images (N = $5$). As D2 pointed out, “Medical images often have a lot of features of different lesions. Even for us, it takes a high level of expertise and careful attention, so I guess it’s a real challenge for AI.” Additionally, many physicians highlighted the issue of imbalanced label distribution in medical images (N = $3$). According to D1, “…a lot of symptoms are rare, and we might only see a few cases over several months, making it hard to capture them during [data] collection.” Furthermore, several physicians discussed the phenomenon of Label Co-occurrence (N = $3$). D4 commented, “Different lesions can be connected or show up together, like certain complications, especially in tongue images, which we might miss if we don’t examine [them] carefully.” These insights underscore the significance of Imbalanced Distribution and Label Co-occurrence as prevalent and critical issues in medical imaging. This leads to the first challenge: C1. The intricate and multifaceted nature of multi-label medical images.
@@ 3.2.2. Challenges in Human-AI Collaboration
Both physicians and engineers acknowledge the benefits of incorporating human knowledge into the model training process, yet they recognize significant challenges (N = $8$). Physicians highlighted that differences in terminology and understanding create communication barriers with engineers, a concern frequently supported by previous studies (Martín-Noguerol et al., 2021), which ultimately impacts the model’s performance. D1 explained, “When I talk with collaborators, I try to simplify the medical terms, but they still often don’t understand or misinterpret what I’m saying, which means we don’t always get the AI results we’re aiming for.” Engineers echoed this sentiment, with P2 stating, “…these terms are new to us, so it’s harder to grasp. We often struggle to fully understand what the physicians mean, which ends up affecting the functionality or outcomes they’re looking for.P3 further noted, “Sometimes, a lot of the medical concepts or rules that physicians suggest are tough to translate into mathematical models during the process.” Consequently, the medical knowledge or rules that physicians aim to integrate into the model are not always effectively communicated or implemented by engineers. This highlights the third challenge: C3. The ineffectiveness and difficulty of integrating human knowledge into model training.
@@
Several physicians have reported a decrease in their trust in AI during collaboration (N = $3$). This issue largely stems from the lack of model interpretability, a concern also highlighted in previous research (A. and R., 2023). As D3 remarked, “I don’t get how the model comes to these conclusions, and sometimes I can’t be sure its decisions match our clinical judgment.” Additionally, mistrust can arise from physicians not being involved in the model training and decision-making processes. D2 pointed out, “Usually, I just get the results from the model, but if I spot [any] errors, I have no way to fix them. This really hurts my trust in the system.” Consequently, the collaboration suffers from inadequate feedback between the parties, which significantly erodes the physicians’ trust. The fourth challenge is: C4. Lack of mutual feedback in the collaboration between physicians and AI models.
@@
When discussing the collaboration between physicians and engineers, both parties raised several specific and nuanced challenges. One major concern was the difference in their working styles (N = $6$). D4 emphasized, “Our work is usually driven by clinical needs and urgency, while they’re more focused on project deadlines and data availability. This difference often causes mismatched timelines and inefficiency.P1 added, “Physicians’ time is incredibly valuable, and they want to see results quickly, but developing and training our models takes time and constant adjustments.” In addition, D5 mentioned, “Previous collaborators gave us some interfaces to help with data analysis or using AI, but I found them hard to use, and some features were tough to understand even after training.” The complexity of existing workflows often demands significant time and effort from physicians to learn and engage in the modeling process, underscoring the fifth challenge: C5. Inconsistency and high learning costs in the collaboration between physicians and engineers.
@@
Furthermore, engineers pointed out that physicians might experience issues related to the “overuse” of AI models (N = $2$). P1 observed, “In a previous project, there was a case where the model picked up incorrect information because physicians used it improperly, so it’s necessary for us to re-evaluate the model’s usability.” Similarly, D5 mentioned, “I’m concerned I might make mistakes without even realizing it, so I prefer to double-check the performance with engineers.” This highlights the sixth challenge: C6. Lack of joint review of the model’s feedback by physicians and engineers.
@@ 3.3. Design Goals
After interviewing both physicians and engineers, we have summarized a series of design goals aimed at effectively addressing the aforementioned challenges. Our approach focuses on improving the traditional collaboration mode between these two parties (Figure 1), enhancing conventional classification methods to foster more direct feedback and interaction between physicians and AI models. We have organized the seven design goals into two key areas: Bias Refinement, which involves Pixel-level Bias Refinement Integrated with Physicians’ Knowledge, and Bias Explanation, which focuses on the Display and Analysis of Biases During AI Training and Fine-tuning Results. These components function within an Iterative Loop, continuously reducing bias, refining the model’s predictive performance, increasing physicians’ engagement and trust, and simultaneously lightening the workload of engineers.
Bias Refinement
@@
Given the differences in professional expertise, work modes, and technical barriers between physicians and engineers, efforts to incorporate medical knowledge into AI training may not always produce optimal results, leading to inefficiencies in collaboration. However, this integration is essential for effective medical image training. To overcome these challenges (C3, C5), we aim to enhance more direct cooperation between physicians and AI models, while also improving physician-engineer collaboration. This brings us to our first design goal: DG1. Facilitate more direct interaction and knowledge integration between physicians and AI models. This objective was emphasized repeatedly during our semi-structured interview.
Bias Explanation
@@
To address the fourth challenge (C4), we should not only facilitate the integration of physicians’ knowledge into AI but also provide meaningful feedback from the AI back to the physicians. This requires clear displays of the model’s training and decision-making processes, enabling physicians to identify biases or inaccuracies generated by the model. For multi-label medical images, it is essential to present the different symptoms within a single image. As D3 emphasized during the semi-structured interview, “I want the model to clearly mark or highlight the locations of each symptom it identifies. I will only trust it when its judgments align with mine.” This underscores the importance of our second design goal: DG2. Provide detailed displays of the model’s decision-making process during training.
Bias Explanation
@@
Beyond providing feedback during AI training, physicians also emphasized the importance of displaying the outcomes of the training process. As D2 mentioned, “I want to see what changes occur after I correct errors and whether these changes improve the model’s performance.” Illustrating how physicians’ interventions affect model accuracy and reduce bias allows them to better understand the impact of their contributions, which in turn helps validate the model’s usability. This brings us to our third design goal: DG3. Display the impact of physicians’ interventions on model outcomes. This goal further supports the feedback mechanisms outlined in C4 and C6.
Bias Explanation
@@
While physicians play a crucial role in verifying the model, the expertise of engineers in evaluating and controlling the model is equally essential. Even when physicians find the model’s interpretability satisfactory, risks such as overfitting [^1] Overfitting happens when training data details are learned too specifically, so performance drops on new data. or gradient vanishing [^2] Gradient vanishing occurs when gradients become very small, slowing or stopping learning. could still arise due to excessive bias refinement or inherent model issues. Identifying and addressing these potential problems to keep the model usable requires engineers’ specialized knowledge. Therefore, it is vital for both physicians and engineers to collaboratively validate the model. To address C6 and supplement DG3, our fourth design goal is: DG4. Provide feedback on model performance to engineers.
Bias Refinement
@@
To mitigate issues of Imbalanced Distribution and Label Co-occurrence in MLMIC (C1), engineers typically rely on mathematical and algorithmic techniques. As P2 noted, “The model’s inaccuracies often stem from focusing on incorrect areas.” By enabling physicians to correct biases that arise during model training before optimization is passed on to engineers, the model’s performance and effectiveness can be significantly improved. However, for ML-naive users, steering AI models can be challenging, as highlighted in C5. Therefore, it’s essential to first provide physicians with an understanding of the data and then offer intuitive, user-friendly tools that enable them to easily identify and correct biases. Our fifth design goal is: DG5. Provide physicians with data cues and accessible tools to adjust and correct biases during model training.
Bias Refinement
@@
The similarity of features and the spatial proximity or overlap in medical images pose significant challenges for effective training. Engineers have observed that for a model to accurately learn the characteristics of various symptoms, it requires more detailed image analysis (C2). However, the annotations needed for this level of detail depend heavily on the expertise of physicians, and such annotated data is often lacking. To address this, we can involve physicians in the annotation process during training, allowing them to highlight key features in the images where detailed learning is required. Thus, our sixth design goal is: DG6. Incorporate techniques that enable the model to discern and learn subtle features in medical images.
Iterative Loop
@@
Our final design goal is: DG7. Develop an interactive interface that facilitates iterative feedback between physicians and AI models. This goal is crafted to support feedback exchange (C4) and foster more direct and closer interactions (C3) between physicians and AI models. Within this iterative loop, the process of physicians providing feedback to the AI and receiving feedback from the AI is seamlessly integrated. Through continuous interaction, physicians can incrementally reduce model biases and enhance their performance. The importance of this goal lies in its potential to gradually increase physicians’ familiarity with human-AI collaboration, allowing them to progressively refine biases and improve the model’s performance.
@@ 3.4. Content Analysis
To develop a user-friendly and accessible annotation method (DG5), we first conducted a comprehensive review of the annotation tools commonly employed by our target users. This review included mainstream annotation tools identified through an extensive search across online platforms such as YouTube and GitHub, recent peer-reviewed literature, and interviews with medical imaging experts. As a result, we identified $5$ annotation methods frequently employed in medical AI annotation workflows. We then systematically analyzed their strengths and limitations (Table 9, Appendix A), evaluating them based on usability, annotation granularity, and their suitability for medical applications.
@@ 4. MEDebiaser
Figure 2. A The Panel View provides components for uploading datasets, selecting models, and setting training parameters. B The Label View includes a table displaying label distribution and a chord diagram showing co-occurrence. C The Attention View displays local explanations for the selected labels. D The Modification View includes
D
1
an Editing Area for fine-tuning attention,
D
2
a Recommendation Area for sorting based on metrics, and
D
3
a Note Area for record-keeping. E The Performance View contains
E
1
a Metrics Area to show the model’s metrics for each round of fine-tuning on specific labels, as well as
E
2
a Comparison Area to display comparisons of local explanations after each round of fine-tuning.
@@
MEDebiaser features five main components: the Panel View (Figure 2- A) allows users to upload datasets, select models, and adjust training parameters (DG1); the Label View (Figure 2- B) displays the distribution of labels in a table and their co-occurrence (DG5); the Attention View (Figure 2- C,   Bias Explanation ) provides local explanations during the training and fine-tuning process (DG2, DG7); the Modification View (Figure 2- D,   Bias Modification ) offers intuitive tools for pixel-level fine-tuning (DG1, DG5, DG6, DG7); and the Performance View (Figure 2- E,   Bias Explanation ) displays prediction results and the impacts of user interventions on model outcomes (DG3, DG4). Users can then initiate a new round of fine-tuning based on the retrained model, following an Iterative Loop process. The MEDebiaser workflow (Figure 3) is structured into three main stages: Loading Dataset and Model, Observing and Modifying Attention, and Evaluating Model Performance. By iteratively executing these stages, MEDebiaser enables physicians to continuously correct biases that arise during model training, seamlessly incorporating their medical expertise into the training process. To be noted, the actual application of the system is not restricted to the specific dataset and model showcased.

### v3

@@ Abstract.
copyright: acmlicensedjournalyear: 2025copyright: ccconference: The 38th Annual ACM Symposium on User Interface Software and Technology; September 28-October 1, 2025; Busan, Republic of Koreabooktitle: The 38th Annual ACM Symposium on User Interface Software and Technology (UIST ’25), September 28-October 1, 2025, Busan, Republic of Koreadoi: 10.1145/3746059.3747725isbn: 979-8-4007-2037-6/2025/09submissionid: 8158ccs: Human-centered computing Human computer interaction (HCI)
@@ 3.1.3. Analysis
We transcribed the audio recordings into text scripts and corrected transcription errors. Using thematic analysis (Oktay, 2012), we conducted a comprehensive examination of the interview data. Initially, all authors read through the scripts to achieve a shared understanding, then proceeded with the coding process. During the initial coding phase, two authors segmented the text data into meaningful units and assigned labels to each segment. Subsequently, two other authors refined and adjusted these labels, uncovering additional patterns and themes. Finally, all authors shared and discussed the coding results, reaching a consensus to ensure consistency and reliability. Based on the coding results, we summarized the codebook (Table 2), which includes 444 themes and 131313 labels, with definitions clarified based on the interviewees’ perspectives. Through this rigorous process, we addressed RQ1 and RQ2, summarizing the challenges and needs faced by physicians and engineers in MLMIC during their collaboration with AI.
@@ 3.2.1. Challenges in Multi-label Medical Image Classification
Physician’s Perspective. All physicians underscored the unique complexities of multi-label medical images (N = 555). As D2 pointed out, “Medical images often have a lot of features of different lesions. Even for us, it takes a high level of expertise and careful attention, so I guess it’s a real challenge for AI.” Additionally, many physicians highlighted the issue of imbalanced label distribution in medical images (N = 333). According to D1, “…a lot of symptoms are rare, and we might only see a few cases over several months, making it hard to capture them during [data] collection.” Furthermore, several physicians discussed the phenomenon of Label Co-occurrence (N = 333). D4 commented, “Different lesions can be connected or show up together, like certain complications, especially in tongue images, which we might miss if we don’t examine [them] carefully.” These insights underscore the significance of Imbalanced Distribution and Label Co-occurrence as prevalent and critical issues in medical imaging. This leads to the first challenge: C1. The intricate and multifaceted nature of multi-label medical images.
@@ 3.2.2. Challenges in Human-AI Collaboration
Both physicians and engineers acknowledge the benefits of incorporating human knowledge into the model training process, yet they recognize significant challenges (N = 888). Physicians highlighted that differences in terminology and understanding create communication barriers with engineers, a concern frequently supported by previous studies (Martín-Noguerol et al., 2021), which ultimately impacts the model’s performance. D1 explained, “When I talk with collaborators, I try to simplify the medical terms, but they still often don’t understand or misinterpret what I’m saying, which means we don’t always get the AI results we’re aiming for.” Engineers echoed this sentiment, with P2 stating, “…these terms are new to us, so it’s harder to grasp. We often struggle to fully understand what the physicians mean, which ends up affecting the functionality or outcomes they’re looking for.P3 further noted, “Sometimes, a lot of the medical concepts or rules that physicians suggest are tough to translate into mathematical models during the process.” Consequently, the medical knowledge or rules that physicians aim to integrate into the model are not always effectively communicated or implemented by engineers. This highlights the third challenge: C3. The ineffectiveness and difficulty of integrating human knowledge into model training.
@@
Several physicians have reported a decrease in their trust in AI during collaboration (N = 333). This issue largely stems from the lack of model interpretability, a concern also highlighted in previous research (A. and R., 2023). As D3 remarked, “I don’t get how the model comes to these conclusions, and sometimes I can’t be sure its decisions match our clinical judgment.” Additionally, mistrust can arise from physicians not being involved in the model training and decision-making processes. D2 pointed out, “Usually, I just get the results from the model, but if I spot [any] errors, I have no way to fix them. This really hurts my trust in the system.” Consequently, the collaboration suffers from inadequate feedback between the parties, which significantly erodes the physicians’ trust. The fourth challenge is: C4. Lack of mutual feedback in the collaboration between physicians and AI models.
@@
When discussing the collaboration between physicians and engineers, both parties raised several specific and nuanced challenges. One major concern was the difference in their working styles (N = 666). D4 emphasized, “Our work is usually driven by clinical needs and urgency, while they’re more focused on project deadlines and data availability. This difference often causes mismatched timelines and inefficiency.P1 added, “Physicians’ time is incredibly valuable, and they want to see results quickly, but developing and training our models takes time and constant adjustments.” In addition, D5 mentioned, “Previous collaborators gave us some interfaces to help with data analysis or using AI, but I found them hard to use, and some features were tough to understand even after training.” The complexity of existing workflows often demands significant time and effort from physicians to learn and engage in the modeling process, underscoring the fifth challenge: C5. Inconsistency and high learning costs in the collaboration between physicians and engineers.
@@
Furthermore, engineers pointed out that physicians might experience issues related to the “overuse” of AI models (N = 222). P1 observed, “In a previous project, there was a case where the model picked up incorrect information because physicians used it improperly, so it’s necessary for us to re-evaluate the model’s usability.” Similarly, D5 mentioned, “I’m concerned I might make mistakes without even realizing it, so I prefer to double-check the performance with engineers.” This highlights the sixth challenge: C6. Lack of joint review of the model’s feedback by physicians and engineers.
@@ 3.3. Design Goals
After interviewing both physicians and engineers, we have summarized a series of design goals aimed at effectively addressing the aforementioned challenges. Our approach focuses on improving the traditional collaboration mode between these two parties (Figure 1), enhancing conventional classification methods to foster more direct feedback and interaction between physicians and AI models. We have organized the seven design goals into two key areas: [Uncaptioned image], which involves Pixel-level Bias Refinement Integrated with Physicians’ Knowledge, and [Uncaptioned image], which focuses on the Display and Analysis of Biases During AI Training and Fine-tuning Results. These components function within an [Uncaptioned image], continuously reducing bias, refining the model’s predictive performance, increasing physicians’ engagement and trust, and simultaneously lightening the workload of engineers.
@@
[Uncaptioned image] Given the differences in professional expertise, work modes, and technical barriers between physicians and engineers, efforts to incorporate medical knowledge into AI training may not always produce optimal results, leading to inefficiencies in collaboration. However, this integration is essential for effective medical image training. To overcome these challenges (C3, C5), we aim to enhance more direct cooperation between physicians and AI models, while also improving physician-engineer collaboration. This brings us to our first design goal: DG1. Facilitate more direct interaction and knowledge integration between physicians and AI models. This objective was emphasized repeatedly during our semi-structured interview.
@@
[Uncaptioned image] To address the fourth challenge (C4), we should not only facilitate the integration of physicians’ knowledge into AI but also provide meaningful feedback from the AI back to the physicians. This requires clear displays of the model’s training and decision-making processes, enabling physicians to identify biases or inaccuracies generated by the model. For multi-label medical images, it is essential to present the different symptoms within a single image. As D3 emphasized during the semi-structured interview, “I want the model to clearly mark or highlight the locations of each symptom it identifies. I will only trust it when its judgments align with mine.” This underscores the importance of our second design goal: DG2. Provide detailed displays of the model’s decision-making process during training.
@@
[Uncaptioned image] Beyond providing feedback during AI training, physicians also emphasized the importance of displaying the outcomes of the training process. As D2 mentioned, “I want to see what changes occur after I correct errors and whether these changes improve the model’s performance.” Illustrating how physicians’ interventions affect model accuracy and reduce bias allows them to better understand the impact of their contributions, which in turn helps validate the model’s usability. This brings us to our third design goal: DG3. Display the impact of physicians’ interventions on model outcomes. This goal further supports the feedback mechanisms outlined in C4 and C6.
@@
[Uncaptioned image] While physicians play a crucial role in verifying the model, the expertise of engineers in evaluating and controlling the model is equally essential. Even when physicians find the model’s interpretability satisfactory, risks such as overfitting [^1] Overfitting happens when training data details are learned too specifically, so performance drops on new data. or gradient vanishing [^2] Gradient vanishing occurs when gradients become very small, slowing or stopping learning. could still arise due to excessive bias refinement or inherent model issues. Identifying and addressing these potential problems to keep the model usable requires engineers’ specialized knowledge. Therefore, it is vital for both physicians and engineers to collaboratively validate the model. To address C6 and supplement DG3, our fourth design goal is: DG4. Provide feedback on model performance to engineers.
@@
[Uncaptioned image] To mitigate issues of Imbalanced Distribution and Label Co-occurrence in MLMIC (C1), engineers typically rely on mathematical and algorithmic techniques. As P2 noted, “The model’s inaccuracies often stem from focusing on incorrect areas.” By enabling physicians to correct biases that arise during model training before optimization is passed on to engineers, the model’s performance and effectiveness can be significantly improved. However, for ML-naive users, steering AI models can be challenging, as highlighted in C5. Therefore, it’s essential to first provide physicians with an understanding of the data and then offer intuitive, user-friendly tools that enable them to easily identify and correct biases. Our fifth design goal is: DG5. Provide physicians with data cues and accessible tools to adjust and correct biases during model training.
@@
[Uncaptioned image] The similarity of features and the spatial proximity or overlap in medical images pose significant challenges for effective training. Engineers have observed that for a model to accurately learn the characteristics of various symptoms, it requires more detailed image analysis (C2). However, the annotations needed for this level of detail depend heavily on the expertise of physicians, and such annotated data is often lacking. To address this, we can involve physicians in the annotation process during training, allowing them to highlight key features in the images where detailed learning is required. Thus, our sixth design goal is: DG6. Incorporate techniques that enable the model to discern and learn subtle features in medical images.
@@
[Uncaptioned image] Our final design goal is: DG7. Develop an interactive interface that facilitates iterative feedback between physicians and AI models. This goal is crafted to support feedback exchange (C4) and foster more direct and closer interactions (C3) between physicians and AI models. Within this iterative loop, the process of physicians providing feedback to the AI and receiving feedback from the AI is seamlessly integrated. Through continuous interaction, physicians can incrementally reduce model biases and enhance their performance. The importance of this goal lies in its potential to gradually increase physicians’ familiarity with human-AI collaboration, allowing them to progressively refine biases and improve the model’s performance.
@@ 3.4. Content Analysis
To develop a user-friendly and accessible annotation method (DG5), we first conducted a comprehensive review of the annotation tools commonly employed by our target users. This review included mainstream annotation tools identified through an extensive search across online platforms such as YouTube and GitHub, recent peer-reviewed literature, and interviews with medical imaging experts. As a result, we identified 555 annotation methods frequently employed in medical AI annotation workflows. We then systematically analyzed their strengths and limitations (Table 9, Appendix A), evaluating them based on usability, annotation granularity, and their suitability for medical applications.
@@ 4. MEDebiaser
Figure 2. Refer to caption The Panel View provides components for uploading datasets, selecting models, and setting training parameters. Refer to caption The Label View includes a table displaying label distribution and a chord diagram showing co-occurrence. Refer to caption The Attention View displays local explanations for the selected labels. Refer to caption The Modification View includes Refer to caption an Editing Area for fine-tuning attention, Refer to caption a Recommendation Area for sorting based on metrics, and Refer to caption a Note Area for record-keeping. Refer to caption The Performance View contains Refer to caption a Metrics Area to show the model’s metrics for each round of fine-tuning on specific labels, as well as Refer to caption a Comparison Area to display comparisons of local explanations after each round of fine-tuning.
@@
MEDebiaser features five main components: the Panel View (Figure 2-[Uncaptioned image]) allows users to upload datasets, select models, and adjust training parameters (DG1); the Label View (Figure 2-[Uncaptioned image]) displays the distribution of labels in a table and their co-occurrence (DG5); the Attention View (Figure 2-[Uncaptioned image],  [Uncaptioned image] ) provides local explanations during the training and fine-tuning process (DG2, DG7); the Modification View (Figure 2-[Uncaptioned image],  [Uncaptioned image] ) offers intuitive tools for pixel-level fine-tuning (DG1, DG5, DG6, DG7); and the Performance View (Figure 2-[Uncaptioned image],  [Uncaptioned image] ) displays prediction results and the impacts of user interventions on model outcomes (DG3, DG4). Users can then initiate a new round of fine-tuning based on the retrained model, following an [Uncaptioned image] process. The MEDebiaser workflow (Figure 3) is structured into three main stages: Loading Dataset and Model, Observing and Modifying Attention, and Evaluating Model Performance. By iteratively executing these stages, MEDebiaser enables physicians to continuously correct biases that arise during model training, seamlessly incorporating their medical expertise into the training process. To be noted, the actual application of the system is not restricted to the specific dataset and model showcased.